The supply function is Q 60 2p - 20r, and the demand function is Q- 260-2p, where r is the rental cost of capital. How do the equilibrium price and quantity vary with r? The effect of r on the equilibrium price is (Enter your resonse as a whole number.) The effect of r on the equilibrium quantity isEnter your resonse as a whole number.) op
Answer #1

According to the question,

The equilibrium in the market will be at the point where quantity demanded is equal to quantity supplied.

Given :

Qd = 80+2p-20r

Qs= 260-2p

Qd =  Qs

80+2p-20r = 260-2p

Reaaranging the terms,

2p+2p = 260-80+20r

4p = 180+20r

Taking 4 common in above equation, we get

p = 45 +5r

Using equation of p in equations Qd or Qs to get equilibrium quantity:

Qe = 80+2*(45 +5r)-20r

Qe= 80+90+10r-20r

Equilibrium Q = 170-10r

Similarly,

Qe = 260-2*(45 +5r)

Qe = 260-90-10r

Qe = 170-10r

Equilibrium price : p = 45 +5r

Equilibrium Quantity: Q = 170-10r

The equilibrium price increases with r as shown in the table below and also the effect of r on equilibrium price is positive (5>0). However, the equilibrium quantity decreases with r as shown in the table below and also the effect of r on equilibrium quantity is negative (-10<0).

r=0

r=1

r=2

r=3

r=4

r=5

p = 45 +5r

45

50

55

60

65

70

Q = 170-10r

170

160

150

140

130

120

The effect of r on equilibrium price is

p = 45 +5r

Differentiating p with respect to r:

dp=5 dr

The effect of r on equilibrium Quantity is

Q = 170-10r

Differentiating Q with respect to r:

de =-10